# Fact :Teaching MLLMs with Faithful, Concise and Transferable Rationales

### v1

@@ Abstract.
The remarkable performance of Multimodal Large Language Models (MLLMs) has unequivocally demonstrated their proficient understanding capabilities in handling a wide array of visual tasks. Nevertheless, the opaque nature of their black-box reasoning processes persists as an enigma, rendering them uninterpretable and struggling with hallucination. Their ability to execute intricate compositional reasoning tasks is also constrained, culminating in a stagnation of learning progression for these models. In this work, we introduce Fact, a novel paradigm designed to generate multimodal rationales that are faithful, concise, and transferable for teaching MLLMs. This paradigm utilizes verifiable visual programming to generate executable code guaranteeing faithfulness and precision. Subsequently, through a series of operations including pruning, merging, and bridging, the rationale enhances its conciseness. Furthermore, we filter rationales that can be transferred to end-to-end paradigms from programming paradigms to guarantee transferability. Empirical evidence from experiments demonstrates the superiority of our method across models of varying parameter sizes, significantly enhancing their compositional reasoning and generalization ability. Our approach also reduces hallucinations owing to its high correlation between images and text.
@@ 1. Introduction
Multimodal Large Language Models (MLLMs) (Dai et al., 2023; Ye et al., 2023, 2024; Chen et al., 2023; Li et al., 2023d, c) enhance the natural and sophisticated interaction between humans and machines, offering distinct advantages in solving visual tasks such as image grounding (Lin et al., 2014), Visual Question Answering (VQA) (Hudson and Manning, 2019; Marino et al., 2019), and scene graph generation (Xu et al., 2017). However, these end-to-end large models exhibit an almost black-box level of interpretability: their reasoning processes remain inexplicable, merely leveraging their formidable representational capabilities to fit spurious correlations (Mitra et al., 2023), thus increasing the risk of hallucinations. Furthermore, the compositional reasoning ability of these models is limited (Ma et al., 2023; Cascante-Bonilla et al., 2023; Doveh et al., 2023) (Figure 1). Even the most advanced state-of-the-art (SOTA) proprietary MLLM, such as GPT-4V (OpenAI et al., 2024), also struggles to perform well in tasks involving counting (Acharya et al., 2018), spatial reasoning (Liu et al., 2023a) and intricate VQA tasks (Marino et al., 2019; Hudson and Manning, 2019).
In this paper, our objective is to enhance the explicit intermediate reasoning capabilities of MLLMs by distilling interpretable rationales. An ideal rationale should embody three properties: 1) Faithfulness. Faithfulness necessitates that the rationale’s reasoning process remains staunchly aligned with the model’s conclusions. This alignment is paramount as it ensures the integrity and reliability of the reasoning process, ultimately leading to conclusions that are both logical and accurate. However, verifying the faithfulness of a rationale poses significant challenges. It is not uncommon for rationales to resemble logical reasoning superficially (Pan et al., 2023), yet upon closer inspection, they may lack a direct correlation to the intended conclusion (Turpin et al., 2023), undermining the model’s effectiveness and trustworthiness. 2) Conciseness. The principle of conciseness stresses the importance of eliminating superfluous in-context information that does not contribute to the reasoning process. The presence of irrelevant information can obfuscate the model’s reasoning pathway, leading to reduced decision-making accuracy (Shi et al., 2023; Jia and Liang, 2017). Extracting a concise and clear rationale is still a process fraught with difficulty, yet such brevity aids in efficiency. 3) Transferability. Transferability refers to the ability of a rationale to encapsulate the model’s explicit inductive understanding of knowledge in a manner that is applicable across various models, paradigms, and scales. A transferable rationale enhances the generalizability of the model’s learned reasoning (Marasović et al., 2022), facilitating the sharing of insights and improvements across different models. However, achieving a high degree of transferability is challenging due to the diverse architectures and learning mechanisms employed by different models.
@@
In this work, we explore a different novel method: Fact, as illustrated in Figure 1, leveraging the powerful symbolic reasoning capabilities of code-pretrained models to construct a faithful, concise, and transferable natural language rationale, which is then applied in teaching MLLMs. Specifically, 1) we generate faithful code by employing a code generation model (OpenAI, [n. d.]) to utilize its compositional capabilities of various tools for the visual task. Then we record the execute trace as a draft chain-of-thought (CoT), retaining only those snippets that compile successfully and yield correct outcomes. 2) We define three operations on the rationales: pruning, merging, and bridging, to simplify code execute trace into natural language by pruning irrelevance in abstract syntax tree (AST), merging duplicates in symbolic traces, and bridging logical gaps to form coherent CoT. A language model (OpenAI, [n. d.]) is then utilized to seamlessly bridge these gaps and refine the narrative. 3) We filter transferable rationales that successfully transfer programmatic reasonings to end-to-end models (Awadalla et al., 2023; Zhu et al., 2023a) and 4) distillate the rationale step-by-step (Hsieh et al., 2023; Park et al., 2023), aiming to enhance the quality and applicability of distilled knowledge across various domains.
In our paradigm, we first ensure that the rationales are correct and rigorous: Given that programs can accurately guide the derivation of answers, the rationales based on these programs can also faithfully represent both the reasoning process and its outcomes (Surís et al., 2023a). Secondly, our rationales are concise and brief: the unexecuted branches within programs are pruned, the repetitive assignments in loops are merged, the repeated uses of tool combinations are inductively summarized, and the gaps in the reasoning process are bridged. Thirdly, our rationales are transferable and consistent: the CoT rationales derived from generated from programming execution traces are selectively retained, only emphasizing those that aid end-to-end models in achieving the correct answers through logical reasoning.
Our experiments demonstrate that CoT rationales, characterized by faithfulness, conciseness, and transferability, generally enhance MLLM performance on downstream tasks (Lin et al., 2014; Hudson and Manning, 2019; Marino et al., 2019; Acharya et al., 2018). Trained with these rationales, MLLMs not only show robust capabilities in counting and compositional reasoning tasks but also guide large vision models to comprehend and reason more logically. The directness of the rationales, devoid of extraneous irrelevance, effectively reduces the occurrence of hallucinations. Ablation studies within our research framework further substantiate the indispensability of these three attributes, affirming the superiority of our method from multiple perspectives.
@@
  • We introduce an innovative paradigm, Fact, to distill code-pretrained models’ logic into a faithful, concise, and transferable rationale for teaching MLLMs, enabling them to harness advanced reasoning capabilities.
  • We refine CoT rationales through controllable editing operations—pruning, merging, and bridging—to enhance their quality and coherence. Furthermore, we have also trained a compact model to identify logical gaps. These operations effectively address challenges in maintaining relevance in distilled knowledge.
  • We validate that the CoT rationale generated through a programming-based approach is applicable for distillation into end-to-end models. It underscores the practicality of our CoT rationales generation paradigm, showcasing its versatility and effectiveness in enhancing model learning.
  • Highlighting the versatility and transferability of our approach, we demonstrate its capability to enhance the reasoning abilities of MLLMs across various sizes. Our evaluation on tasks like GQA (Hudson and Manning, 2019), OKVQA (Marino et al., 2019), TallyQA (Acharya et al., 2018), and COCO (Lin et al., 2014) reveals significant performance improvements in MLLMs, and effectively reduces hallucinations by being devoid of extraneous irrelevance.
@@ 2.1. Visual Program Distillation
Visual programming is a burgeoning field that employs neural symbols (Gupta and Kembhavi, 2023) or Python modules (Surís et al., 2023b) for task synthesis and execution. While it offers enhanced performance and interpretability through precise image manipulation via code, its dependency on multiple models and prolonged inference times necessitates substantial computational resources. In contrast, our approach and Visual Program Distillation (VPD) (Hu et al., 2023) diverge significantly in handling program-based methods. VPD simplifies multimodal learning by distilling tool use and programmatic reasoning into smaller models but retains unnecessary execution traces, lacks precise spatial task execution, and overlooks the verification of transferability. Our method addresses these limitations through Efficient Rationale Editing: Unlike VPD, we implement controllable editing to refine programs into concise rationales; Enhanced Spatial Task Execution: Our approach exhibits superior spatial reasoning capabilities, enabling more accurate completion of spatial tasks compared to VPD; Verified Transferability: We also emphasize and validate the transferability of program-based rationales to end-to-end models which VPD simply ignored it. These distinctions underscore our contribution to integrating complex reasoning within MLLMs.
@@ 2.2. Rationale Distillation
Deploying LLMs in practical applications is challenging due to their substantial memory and computational demands. A feasible approach to mitigate these challenges is training smaller, task-specific models via fine-tuning or distillation with labels generated by LLMs. While effective, these strategies often necessitate extensive training data to match the performance with LLMs (Liu et al., 2023c, b). As a solution, Distillation Step-by-Step (Hsieh et al., 2023) emerges, enabling smaller models to potentially surpass LLMs with less training data. However, existing efforts rarely account for the quality of distillation data and its transferability between large and small models. This neglect can result in the distillation process incorporating irrelevant information, which diverts the smaller model and detracts from its performance. Our approach emphasizes the quality and relevance of distilled information, tackling the common pitfalls of distillation. Furthermore, our method stands out by offering verifiability and transferability. This ensures that the distilled knowledge is not only accurate and relevant but also adaptable across different models and tasks. The ability to verify the correctness and applicability of our CoT rationales sets our work apart, underscoring its novelty and effectiveness in the context of distillation methodologies.
@@ 3.1. Faithful Program Generation
Figure 4. We show several examples of the process that generates CoT rationale for distillation.
@@ 3.2. Concise CoT Conversion
Directly utilizing code for distillation learning is not advantageous due to redundancies within the code, like unmet conditionals in if statements, and repetitive outputs from for loops. Previous work (Shi et al., 2023; Jia and Liang, 2017) demonstrated that models are prone to distraction by irrelevance. Moreover, for models not trained on code corpora, understanding programming becomes significantly challenging. In light of this, we propose three operations to transform program traces into concise CoT rationales: dynamic pruning, symbolic merging, and logical bridging. These operations aim to refine the generated CoTs by removing irrelevance, combining repetition for conciseness, and bridging gaps in logicen for coherence, thereby enhancing the quality and relevance of the distilled content for MLLMs.
Dynamic Pruning: The reason for taking the pruning operation is that the rationale should explain why it is, rather than why it is not. To this end, we initially construct the code $z$ into an AST, denoted $t=\{V,E\}$, where $V=\{v_{1},...,v_{n}\}$ represents the vertices and $E=\{e_{1},...,e_{m}\}$ represents the edges. The AST facilitates more granular modifications via node editing compared to direct manipulations of the code text alone. Given a specific image, corresponding query, and the generated program, the final output answer will be unique. Thus, variables and code statements that are not used in the execution process should be discarded. For example, within conditional statements (if), if the condition is not met, we should not record this condition as false, since the program does not enter the scope of the if statement. By doing this, we can remove content that is irrelevant to the context of the reasoning process.
@@ 3.3. CoT Transferability Verification
3.4. Distillation Step-by-Step
$L_{label}=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i},q_{i}),\hat{y}_{i})$
$L_{rationale}=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i},q_{i}),\hat{r}_{i})$
$L=L_{label}+\lambda L_{rationale}$
Figure 5. Example outputs of MiniGPT4 trained with Fact.
@@ 4.1. Experimental Setup
Table 2. Comparison of MME and POPE benchmark.
MME POPE
@@
OpenFlamingo (Awadalla et al., 2023) 668.2 52.6 67.2 56.0
   w/ Instruct generalist 847.3 69.5 73.1 68.4
   w/ Fact generalist 912.2 73.0 75.6 71.5
MiniGPT4 (Zhu et al., 2023a) 581.7 43.3 50.8 47.9
   w/ Instruct generalist 864.9 68.3 74.4 71.2
   w/ Fact generalist 1034.7 78.7 83.7 79.1
mPlug-Owl (13B) pre-trained GQA OK-VQA
   + in-context prompt 56.5 57.6
   + CoT (OpenFlamingo-Fact) 62.7 62.6
   + CoT (MiniGPT4-Fact) 63.2 63.0
@@
Benchmark. Our models are rigorously evaluated on a comprehensive suite of zero-shot benchmarks to ascertain their performance across various tasks. Specifically, for image captioning, we utilized datasets from COCO (Lin et al., 2014) and Flickr 30K (Plummer et al., 2015), with the model’s performance assessed using the CIDEr scoring metric. For VQA, we employed the VQAv2 (Goyal et al., 2017) dataset, evaluating the model on the test-dev split. The GQA (Hudson and Manning, 2019) benchmark was assessed on its test-dev set, while the OK-VQA (Marino et al., 2019) was evaluated on the val split. For counting tasks, we used the TallyQA (Acharya et al., 2018) benchmark, assessing our model on the test set. In addition to these primary benchmarks, we extended our evaluation to include additional benchmarks such as the MME (Fu et al., 2024) score and POPE (Li et al., 2023a) F1-score, which are designed to further test the model’s capabilities in multi-modal understanding and predictive object positioning, respectively. We present the prompts required for evaluating downstream tasks in Appendix B.

### v2

@@ Abstract.
The remarkable performance of Multimodal Large Language Models (MLLMs) has demonstrated their proficient understanding capabilities in handling various visual tasks. Nevertheless, the opaque nature of black-box reasoning processes persists as an enigma, rendering them uninterpretable and struggling with hallucination. Their ability to execute intricate reasoning tasks is also constrained, culminating in stagnation of progression. In this work, we introduce Fact, a novel paradigm designed to generate multimodal rationales that are faithful, concise, and transferable for teaching MLLMs. This paradigm utilizes verifiable visual programming to generate executable code guaranteeing faithfulness. Through a series of operations including pruning, merging, and bridging, the rationale enhances its conciseness. Furthermore, we filter rationales that can be transferred to end-to-end paradigms from programming paradigms to guarantee transferability. Empirical evidence from experiments demonstrates the superiority of Fact across models of varying parameter sizes, significantly enhancing their compositional reasoning and generalization ability and reducing hallucinations owing to its high correlation between images and text.
Visual Programming, Multimodel Chain-of-Thought, Distillation Step-by-Step
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681552isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision problems
@@ 1. Introduction
Multimodal Large Language Models (MLLMs) (Dai et al., 2023; Chen et al., 2023; Li et al., 2023e, c; Wu et al., 2024; Ge et al., 2024; Li et al., 2024) enhance the natural and sophisticated interaction between humans and machines, offering distinct advantages in alignment (Li et al., 2022) and solving visual tasks such as image grounding (Lin et al., 2014), Visual Question Answering (VQA) (Hudson and Manning, 2019; Marino et al., 2019), and scene graph generation (Xu et al., 2017; Fei et al., 2023). However, these end-to-end large models exhibit an almost black-box level of interpretability: their reasoning processes remain inexplicable, merely leveraging their formidable representational capabilities to fit spurious correlations (Mitra et al., 2023), thus increasing the risk of hallucinations. Furthermore, the compositional reasoning ability of these models is limited (Ma et al., 2023; Cascante-Bonilla et al., 2023; Doveh et al., 2023) (Figure 1). Even the most advanced state-of-the-art (SOTA) proprietary MLLM, such as GPT-4V (OpenAI et al., 2024), also struggles to perform well in tasks involving counting (Acharya et al., 2018), spatial reasoning (Liu et al., 2023a; Fei et al., 2024b) and intricate VQA tasks (Marino et al., 2019; Hudson and Manning, 2019).
@@
In this paper, we aim to enhance the explicit intermediate reasoning capabilities of MLLMs by distilling interpretable rationales. An ideal rationale should meet three criteria: 1) Faithfulness. The rationale should staunchly align with the model’s conclusions, ensuring its reliability and logical consistency. This alignment is crucial, yet challenging to verify, as rationales may appear logical but lack a true connection to the conclusions, which can diminish trust in the model (Pan et al., 2023; Turpin et al., 2023). 2) Conciseness. The conciseness stresses the importance of eliminating superfluous in-context information that does not contribute to the reasoning process. The irrelevant information can obfuscate the model’s reasoning pathway (Shi et al., 2023; Jia and Liang, 2017). Extracting a concise rationale is still a process fraught with difficulty, yet such brevity aids in efficiency. 3) Transferability. Transferability refers to the ability of a rationale to encapsulate the model’s explicit inductive understanding of knowledge in a manner that is applicable across various models, paradigms, and scales. A transferable rationale enhances the generalizability of the model’s learned reasoning (Marasović et al., 2022), facilitating the sharing of insights across different models.
@@
In this work, we explore a novel method: Fact, as illustrated in Figure 1, leveraging the symbolic reasoning capabilities of code-pretrained models to construct a faithful, concise, and transferable rationale, which is then applied in teaching MLLMs. Specifically, 1) we generate faithful code by a code generation model (OpenAI, [n. d.]) to utilize its compositional capabilities for the visual task. Then we record the execute trace as a draft chain-of-thought (CoT), retaining snippets that compile successfully and yield correct outcomes. 2) We define three operations: pruning, merging, and bridging, to simplify execute trace into natural language by pruning irrelevance in abstract syntax tree (AST), merging duplicates in symbolic traces, and bridging logical gaps to form coherent CoT. A language model (OpenAI, [n. d.]) is then utilized to bridge these gaps and refine the narrative. 3) We filter transferable rationales that successfully transfer programmatic reasonings to end-to-end models (Awadalla et al., 2023; Fei et al., 2024a; Zhu et al., 2023a) and 4) distillate the rationale step-by-step (Hsieh et al., 2023; Park et al., 2023), enhancing the applicability of distilled knowledge across various domains.
In our paradigm, we first ensure the rationales are correct and rigorous: Given that programs can accurately guide the derivation of answers, the rationales based on these programs can faithfully represent both the reasoning process and outcomes (Surís et al., 2023a). Secondly, our rationales are concise and brief: the unexecuted branches are pruned, the repetitive assignments in loops are merged, the repeated uses of tool combinations are inductively summarized, and the gaps in the reasoning process are bridged. Thirdly, our rationales are transferable and consistent: the CoT rationales derived from generated from programming execution traces are selectively retained, only emphasizing those that aid end-to-end models in achieving the correct answers through logical reasoning.
Our experiments demonstrate that CoT rationales, characterized by faithfulness, conciseness, and transferability, enhance MLLM performance on downstream tasks (Lin et al., 2014; Hudson and Manning, 2019; Marino et al., 2019; Acharya et al., 2018). Trained with these rationales, MLLMs show robust capabilities in counting and compositional reasoning tasks and can guide large vision models to comprehend more logically. The directness of the rationales, devoid of extraneous irrelevance, effectively reduces the occurrence of hallucinations. Ablation studies within our research framework further substantiate the indispensability of these three attributes, affirming the superiority of our method from multiple perspectives.
@@
  • We introduce an innovative paradigm, Fact, to distill code-pretrained models’ logic into a faithful, concise, and transferable rationale for teaching MLLMs to advance in reasoning.
  • We refine CoT rationales through controllable editing operations—pruning, merging, and bridging—to enhance their quality. Furthermore, we also train a compact model to identify logical gaps. These operations effectively address challenges in maintaining relevance in distilled knowledge.
  • We validate that the CoT rationale generated through a programming-based approach is applicable for distillation into end-to-end models.
  • Highlighting the versatility and transferability of our approach, we evaluate Fact on visual datasets like GQA (Hudson and Manning, 2019), OKVQA (Marino et al., 2019), TallyQA (Acharya et al., 2018), and COCO (Lin et al., 2014), revealing significant performance improvements and effectively reducing hallucinations by being devoid of extraneous irrelevance.
@@ 2.1. Visual Program Distillation
Visual programming is a burgeoning field that employs neural symbols (Gupta and Kembhavi, 2023) or Python modules (Surís et al., 2023b; Gao et al., 2023) for task synthesis and execution. While it offers enhanced performance and interpretability through precise image manipulation via code, its dependency on multiple models and prolonged inference times necessitates substantial computational resources. In contrast, our approach and Visual Program Distillation (VPD) (Hu et al., 2023) diverge significantly in handling program-based methods. VPD simplifies multimodal learning by distilling tool use and programmatic reasoning into smaller models but retains unnecessary execution traces, lacks precise spatial task execution, and overlooks the verification of transferability. Our method addresses these limitations through Efficient Rationale Editing: Unlike VPD, we implement controllable editing to refine programs into concise rationales; Enhanced Spatial Task Execution: Our approach exhibits superior spatial reasoning capabilities, enabling more accurate completion of spatial tasks compared to VPD; Verified Transferability: We also emphasize and validate the transferability of program-based rationales to end-to-end models which VPD simply ignored it. These distinctions underscore our contribution to integrating complex reasoning within MLLMs.
@@ 2.2. Rationale Distillation
Deploying LLMs in practical applications is challenging due to their substantial memory and computational demands. A feasible approach to mitigate these challenges is training smaller, task-specific models via fine-tuning or distillation with labels generated by LLMs. While effective, these strategies often necessitate extensive training data to match the performance with LLMs (Liu et al., 2023c, b). As a solution, Distillation Step-by-Step (Hsieh et al., 2023) emerges, enabling smaller models to potentially surpass LLMs with less training data. However, existing efforts rarely account for the quality of distillation data and its transferability between large and small models. This neglect can result in the distillation process incorporating irrelevant information. Our approach emphasizes the quality and relevance of distilled information, tackling the common pitfalls of distillation. Furthermore, our method stands out by offering verifiability and transferability. This ensures that the distilled knowledge is not only accurate and relevant but also adaptable across different models and tasks. The ability to verify the correctness and applicability of our CoT rationales sets our work apart, underscoring its novelty and effectiveness in the context of distillation methodologies.
@@ 3.2. Concise CoT Conversion
Directly utilizing code for distillation is not advantageous due to redundancies within the code, like unmet conditionals in if statements, and repetitive outputs from for loops. Previous work (Shi et al., 2023; Jia and Liang, 2017) demonstrated that models are prone to distraction by irrelevance. Moreover, for models not trained in code, understanding programming becomes challenging. In light of this, we propose three operations to transform program traces into concise CoT rationales: dynamic pruning, symbolic merging, and logical bridging. These operations aim to refine the CoTs by removing irrelevance, combining repetition, and bridging gaps in logic, thereby enhancing the quality and relevance of the distilled content for MLLMs.
Dynamic Pruning: The rationale should explain why it is, rather than why it is not. To this end, we construct the code $z$ into an AST, denoted $t=\{V,E\}$, where $V=\{v_{1},...,v_{n}\}$ represents the vertices and $E=\{e_{1},...,e_{m}\}$ represents the edges. The AST facilitates more granular modifications via node editing compared to direct manipulations of the code text. Given a specific image, corresponding query, and the generated program, the final output answer will be unique. Thus, variables and code statements that are not used in the execution process should be discarded. For example, within conditional statements (if), if the condition is not met, we should not record this condition as false, since the program does not enter the scope of the if statement. By doing this, we can remove content that is irrelevant to the context of the reasoning process.
@@
Figure 4. An example of (a) the process that generates CoT rationale for distillation and (b) outputs of MiniGPT4 with Fact.
@@ 3.3. CoT Transferability Verification
3.4. Distillation Step-by-Step
$L_{label}=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i},q_{i}),\hat{y}_{i})$
$L_{rationale}=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i},q_{i}),\hat{r}_{i})$
$L=L_{label}+\lambda L_{rationale}$
@@ 4.1. Experimental Setup
Table 2. Comparison of various benchmarks.
MM-Bench SEED MME POPE
@@
OpenFlamingo (Awadalla et al., 2023) 13.6 33.1 668.2 52.6 67.2 56.0
w/ Instruct generalist 14.8 34.9 847.3 69.5 73.1 68.4
w/ Fact generalist 18.3 39.2 912.2 73.0 75.6 71.5
MiniGPT4 (Zhu et al., 2023a) 32.7 42.8 581.7 43.3 50.8 47.9
w/ Instruct generalist 35.4 44.6 864.9 68.3 74.4 71.2
w/ Fact generalist 40.3 48.5 1034.7 78.7 83.7 79.1
@@
Benchmark. Our models are rigorously evaluated on a comprehensive suite of zero-shot benchmarks to ascertain their performance across various tasks. Specifically, for image captioning, we utilized datasets from COCO (Lin et al., 2014) and Flickr 30K (Plummer et al., 2015), with the model’s performance assessed using the CIDEr scoring metric. For VQA, we employed the VQAv2 (Goyal et al., 2017) dataset, evaluating the model on the test-dev split. The GQA (Hudson and Manning, 2019) benchmark was assessed on its test-dev set, while the OK-VQA (Marino et al., 2019) was evaluated on the val split. For counting tasks, we used the TallyQA (Acharya et al., 2018) benchmark, assessing our model on the test set. In addition to these primary benchmarks, we extended our evaluation to include additional benchmarks such as the MME (Fu et al., 2024), MM-bench (Liu et al., 2024), SEED (Li et al., 2023d) score and POPE (Li et al., 2023a) F1-score, which are designed to further test the model’s capabilities in multi-modal understanding and predictive object positioning, respectively. We present the prompts required for evaluating downstream tasks in Appendix B.